# Free multiderivations of connected subgraph arrangements

### v2

@@ 4.1. Investigating $G_{1}$ and $G_{2}$
The proof in [CK24, §6.2] of the non-freeness of the connected subgraph arrangements stemming from $G_{1}$ up to $G_{8}$ depends on tedious computer calculations. In this section, we provide a non-computational, conceptional proof giving a complete list of free constant multiplicities on the family of graphs from Theorem 2.23, as well as giving a non-computational proof for the non-freeness of ${\mathscr{A}}_{G_{1}}$ and ${\mathscr{A}}_{G_{2}}$ in the process. Combined with Proposition 3.1 we obtain a proof for the non-freeness of the simple connected subgraph arrangements stemming from $G_{1}$ up to $G_{20}$ free of any machine calculations. We use the following method for calculating $\operatorname{LMP}(2)$ of the connected subgraph arrangements. Then we utilize Lemma 2.32.
@@
    For every subgraph $G^{\prime}$ from step $1$ we determine the number of triples $(G^{\prime},G[I],G[J])$, where $G[I],G[J]$ are connected induced subgraphs of $G^{\prime}$ with $V(G[I])\cup V(G[J])=V(G^{\prime})$. Adding up the number of all such triples for all connected induced subgraphs $G^{\prime}$ from (1) provides an invariant $N_{1}$. Using Lemma 2.20 we see that for a fixed graph $G^{\prime}$ this is just the number of rank $2$ localizations with $V(G^{\prime})=A_{i_{3}}$, with the notation as in the lemma.
@@
    Calculate $N_{2}:=\binom{|{\mathscr{A}}_{G}|}{2}$. If all rank $2$ localizations were to consist of exactly two elements, then $N_{2}$ would be the number of all rank $2$ localizations.
@@
    We have $\operatorname{LMP}(2)=N_{2}-N_{1}$. To see this, a rank $2$ localization with two elements is free with exponents $(1,1)$, while a rank $2$ localization with three elements is free with exponents $(1,2)$. So if instead of having three rank $2$ localizations with two elements we have one rank $2$ localization with three elements, then $\operatorname{LMP}(2)$ gets decreased by $1$ (since we assume for $N_{2}$ that all localizations contain exactly two elements). Since $N_{1}$ is the number of rank $2$ localizations with three elements, $N_{2}-N_{1}$ coincides with $\operatorname{LMP}(2)$.
@@
  • $5\times P_{2},2\times P_{3},2\times C_{3},1\times G_{1}$
  • $G^{\prime}=P_{2}:1\times(P_{2},P_{1},P_{1})$
    $G^{\prime}=P_{3}:2\times(P_{3},P_{2},P_{1})$
    $G^{\prime}=C_{3}:3\times(C_{3},P_{2},P_{1})$
    $G^{\prime}=G_{1}:2\times(G_{1},C_{3},P_{1}),2\times(G_{1},P_{3},P_{1}),2\times(G_{1},P_{2},P_{2})$
    $N_{1}=5\cdot 1+2\cdot 2+2\cdot 3+1\cdot 6=21$
  • $N_{2}=\binom{14}{2}=91$
@@
  • $6\times P_{2},4\times C_{3},1\times G_{1}$
  • $G^{\prime}=P_{2}:1\times(P_{2},P_{1},P_{1})$
    $G^{\prime}=C_{3}:3\times(C_{3},P_{2},P_{1})$
    $G^{\prime}=G_{2}:4\times(G_{1},C_{3},P_{1}),3\times(G_{1},P_{2},P_{2})$
    $N_{1}=6\cdot 1+4\cdot 3+1\cdot 7=25$
  • $N_{2}=\binom{15}{2}=105$
@@
Since $|{\mathscr{A}}_{G_{1}}|=14$, a tuple of potential exponents for ${\mathscr{A}}_{G_{1}}$ has to be $(1,4,4,5)_{\leq}$, or less balanced. Calculating $\operatorname{GMP}(2)$ for this sequence gives $\operatorname{GMP}(2)=69$. Since exponents that are less balanced afford a smaller $\operatorname{GMP}(2)$ and since $\operatorname{LMP}(2)=70$, we have $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so ${\mathscr{A}}_{G_{1}}$ is not free, by Theorem 2.28.
Since $|{\mathscr{A}}_{G_{1}}|=15$, choose potential balanced exponents $(1,4,5,5)_{\leq}$ for ${\mathscr{A}}_{G_{2}}$ to maximize $\operatorname{GMP}(2)$. Using this tuple, we get $\operatorname{GMP}(2)=79$. Since exponents that are less balanced yield a smaller value for $\operatorname{GMP}(2)$ and since $\operatorname{LMP}(2)=80$, we have $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so ${\mathscr{A}}_{G_{2}}$ can’t be free, thanks to Theorem 2.28. ∎
@@
Next we modify the algorithm from Remark 4.1 to calculate $\operatorname{LMP}(2)$ when ${\mathscr{A}}_{G}$ gets equipped with a constant multiplicity $\mu=k\mathbbm{1}$ for $k\neq 1$.
@@
Algorithm to calculate $\operatorname{LMP}(2)$ for a constant multiplicity $\mu\equiv k$:
@@
    For every subgraph $G^{\prime}$ from step (1) we determine the number of triples $(G^{\prime},G[I],G[J])$, where $G[I],G[J]$ are connected induced subgraphs of $G^{\prime}$ with $V(G[I])\cup V(G[J])=V(G^{\prime})$. Let $N_{1}$ be the sum of the number of all such triples for all connected induced subgraphs $G^{\prime}$ from part (1).
@@
    Calculate $N_{2}:=\binom{|{\mathscr{A}}_{G}|}{2}$.
@@
    Since $N_{1}$ is the number of rank $2$ localizations with three elements, the number of rank $2$ localizations with two elements is equal to $N_{3}:=N_{2}-3N_{1}$, because if there is one rank $2$ localization with three elements, then the hyperplanes of this localization could have potentially formed three rank $2$ localizations with two elements.
@@
    Using Theorem 2.16 we derive the exponents of all rank $2$ localizations with three elements (which depends on the parity of $k$), while localizations of rank $2$ with two elements are free with exponents $(k,k)$. We get
    $\operatorname{LMP}(2)=\begin{cases}N_{1}\cdot\left(\frac{3k}{2}\right)^{2}+N_{3}k^{2},&\text{ (if }k\text{ is even),}\\ N_{1}\cdot\left(\frac{3k-1}{2}\cdot\frac{3k+1}{2}\right)+N_{3}k^{2},&\text{ (if }k\text{ is odd).}\end{cases}$
@@
For $G\in\{G_{1},G_{2}\}$ and $\mu\equiv k\geq 1$ a constant multiplicity on ${\mathscr{A}}_{G}$, $({\mathscr{A}}_{G},\mu)$ is not free.
@@
The case $k=1$ is Proposition 4.3. So let $k\geq 2$. From the calculation in the example we know that: $G_{1}:$ $N_{1}=21,N_{2}=91$ and $G_{2}:$ $N_{1}=25,N_{2}=105$, which gives $N_{3}=28$ for $G_{1}$ and $N_{3}=30$ for $G_{2}$. So for $G_{1}$ we get
$\operatorname{LMP}(2)=\begin{cases}21\cdot\left(\frac{3k}{2}\right)^{2}+28k^{2}=75.25k^{2},&\text{if $k$ is even.}\\ 21\cdot\left(\frac{3k-1}{2}\cdot\frac{3k+1}{2}\right)+28k^{2}=75.25k^{2}-5.25,&\text{if $k$ is odd,}\end{cases}$
@@
$\operatorname{LMP}(2)=\begin{cases}25\cdot\left(\frac{3k}{2}\right)^{2}+30k^{2}=86.25k^{2},&\text{if $k$ is even.}\\ 25\cdot\left(\frac{3k-1}{2}\cdot\frac{3k+1}{2}\right)+30k^{2})=86.25k^{2}-6.25,&\text{if $k$ is odd.}\end{cases}$
If we assume perfectly balanced exponents for ${\mathscr{A}}_{G_{1}}$ (assuming $14k\equiv 0\mod 4$) we get the set of exponents $(\frac{14k}{4},\frac{14k}{4},\frac{14k}{4},\frac{14k}{4})_{\leq}$ which results in $\operatorname{GMP}(2)=73.5k^{2}$.
If we assume perfectly balanced exponents for ${\mathscr{A}}_{G_{2}}$ (assuming $15k\equiv 0\mod 4$) we get the set of exponents $(\frac{15k}{4},\frac{15k}{4},\frac{15k}{4},\frac{15k}{4})_{\leq}$ which results in $\operatorname{GMP}(2)=84.375k^{2}$.
Since $k\geq 2$, in any case, we have $\operatorname{LMP}(2)>\operatorname{GMP}(2)$ and thus $({\mathscr{A}}_{G_{2}},\mu)$ can’t be free, by Theorem 2.28. ∎
@@ 4.2. Constant multiplicities on a free ${\mathscr{A}}_{G}$
In this subsection, we inspect free constant multiplicities on a free ${\mathscr{A}}_{G}$. We give a complete list when $({\mathscr{A}}_{G},\mu)$ is free for $G\in\{P_{n},A_{n,k},\Delta_{n,k},C_{n}\}$ and $\mu\equiv k\geq 1$.
First we consider the case $G=P_{n}$. Terao showed in [T02] that every constant multiplicity on a Coxeter arrangement is free. Since ${\mathscr{A}}_{P_{n}}$ is the Coxeter arrangement of type $A_{n}$, we derive the freeness of $({\mathscr{A}}_{P_{n}},\mu)$ for $\mu\equiv k\geq 1$.
@@
Let $G=C_{3}$ and $\mu\equiv k\geq 1$ is a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $k\in\{1,3\}$.
@@
For the forward implication, using the algorithm above, we get $N_{1}=1\cdot 3+3\cdot 1=6,N_{2}=\binom{7}{2}=21,N_{3}=21-3\cdot 6=3$. This gives
$\operatorname{LMP}(2)=\begin{cases}6\cdot\left(\frac{3k}{2}\right)^{2}+3k^{2}=16.5k^{2}&\text{ (if }k\text{ is even), or }\\ 6\cdot\left(\frac{3k-1}{2}\cdot\frac{3k+1}{2}\right)+3k^{2}=16.5k^{2}-1.5&\text{ (if }k\text{ is odd).}\end{cases}$
@@
$\operatorname{GMP}(2)=\begin{cases}16\frac{1}{3}k^{2},&\text{if }7k\equiv 0\mod 3\text{ or}\\ 16\frac{1}{3}k^{2}-\frac{1}{3},&\text{else.}\end{cases}$
For $k\geq 4$ we get that $\operatorname{LMP}(2)>\operatorname{GMP}(2)$. For $k=3$ we get equality and for $k=2$ we get $\operatorname{LMP}(2)=66$ and depending on whether the assumed exponents are $(3,5,6)_{\leq},(4,4,6)_{\leq}$, or $(4,5,5)_{\leq}$, a value of $63,64$, or $65$, respectively for $\operatorname{GMP}(2)$. So if $k\not\in\{1,3\}$ we have $\operatorname{LMP}(2)>\operatorname{GMP}(2)$ and thanks to Theorem 2.28, we deduce non-freeness.
@@
Let $G\in\{C_{n},\Delta_{n,k},A_{n,k}\}$. Suppose $\operatorname{rank}{\mathscr{A}}_{G}\geq 4$ and $\mu\equiv k\geq 1$ a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $k=1$.
@@
First let $G\in\{C_{n},\Delta_{n,k}\}$. Since $G$ either has $C_{3}$ as an induced subgraph or becomes $C_{3}$ after contracting edges, we can use Lemma 2.22 and Proposition 4.7 to focus on the cases where $k\in\{1,3\}$. If $k=1$ we know that $({\mathscr{A}}_{G},\mu)$ is free. So let $\mu\equiv 3$ and $G=C_{4}$.
Calculating $\operatorname{LMP}(2)$ we get $N_{1}=1\cdot(4+2)+4\cdot 2+4\cdot 1=18,N_{2}=\binom{13}{2}=78,N_{3}=78-3\cdot 18=24$ and so
@@
Let $G=A_{3,2}$. Let $\mu\equiv k$, then we get
$\operatorname{GMP}(2)=\begin{cases}45.375k^{2},&\text{if }11k\equiv 0\mod 4,\\ 45.375k^{2}-\frac{1}{2},&\text{if }11k\equiv 2\mod 4,\\ 45.375k^{2}-\frac{3}{8},&\text{else.}\\ \end{cases}$
Moreover, $\operatorname{LMP}(2)=46k^{2}$ (if $k$ is even) and $346k^{2}-$ 3 (if $k$ is odd). Once more we see that $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so, by Theorem 2.28, $({\mathscr{A}},\mu)$ fails to be free. Since every $A_{n,k}$ with $n\geq 4$ can be transformed into $A_{3,2}$ after contracting edges, the result follows from Lemma 2.22. ∎
@@
Let $G\in\{P_{n},A_{n,k},\Delta_{n,k},C_{n}\}$ , $\mu\equiv k,(k>1)$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $G=P_{n}$ or $G=C_{3}$ and $k=3$.

### v3

@@ 4.1. Investigating $G_{1}$ and $G_{2}$
The proof in [5, §6.2] of the non-freeness of the connected subgraph arrangements stemming from $G_{1}$ up to $G_{8}$ depends on tedious computer calculations. In this section, we provide a non-computational, conceptional proof giving a complete list of free constant multiplicities on the family of graphs from Theorem 2.23, as well as giving a non-computational proof for the non-freeness of ${\mathscr{A}}_{G_{1}}$ and ${\mathscr{A}}_{G_{2}}$ in the process. Combined with Proposition 3.1 we obtain a proof for the non-freeness of the simple connected subgraph arrangements stemming from $G_{1}$ up to $G_{20}$ free of any machine calculations. We use the following method for calculating $\operatorname{LMP}(2)$ of the connected subgraph arrangements. Then we utilize Lemma 2.32.
@@
    Determine the number of all types (i.e. path-graph, almost-path-graph, etc.) of connected induced subgraphs $G[S]$ of $G=(N,E)$, where $S\subseteq N$ and $|S|\geq 2$.
@@
    For every subgraph $G[S]$ from step $(1)$ we determine the number of triples of the form $(G[S],G[I],G[J])$, where $G[I],G[J]$ are connected induced subgraphs of $G[S]$ with $I\,\dot{\cup}\,J=S$. Adding up the number of all such triples for all connected induced subgraphs $G^{\prime}$ from (1) provides an invariant $\nu_{1}$. Using Lemma 2.20 we see that for a fixed graph $G^{\prime}$ this is just the number of rank $2$ localizations with $N(G^{\prime})=A_{i_{3}}$, with the notation as in the lemma.
@@
    Calculate $\nu_{2}:=\binom{|{\mathscr{A}}_{G}|}{2}$. If all rank $2$ localizations were to consist of exactly two elements, then $\nu_{2}$ would be the number of all rank $2$ localizations.
@@
    We have $\operatorname{LMP}(2)=\nu_{2}-\nu_{1}$. To see this, a rank $2$ localization with two elements is free with exponents $(1,1)$, while a rank $2$ localization with three elements is free with exponents $(1,2)$. So if instead of having three rank $2$ localizations with two elements we have one rank $2$ localization with three elements, then $\operatorname{LMP}(2)$ gets decreased by $1$ (since we assume for $\nu_{2}$ that all localizations contain exactly two elements). Since $\nu_{1}$ is the number of rank $2$ localizations with three elements, $\nu_{2}-\nu_{1}$ coincides with $\operatorname{LMP}(2)$.
@@
  • $5\cdot P_{2},2\cdot P_{3},2\cdot C_{3},1\cdot G_{1}$
  • $G^{\prime}=P_{2}:1\cdot(P_{2},P_{1},P_{1})$
    $G^{\prime}=P_{3}:2\cdot(P_{3},P_{2},P_{1})$
    $G^{\prime}=C_{3}:3\cdot(C_{3},P_{2},P_{1})$
    $G^{\prime}=G_{1}:2\cdot(G_{1},C_{3},P_{1}),2\cdot(G_{1},P_{3},P_{1}),2\cdot(G_{1},P_{2},P_{2})$
    $\nu_{1}=5\cdot 1+2\cdot 2+2\cdot 3+1\cdot 6=21$
  • $\nu_{2}=\binom{14}{2}=91$
@@
  • $6\cdot P_{2},4\cdot C_{3},1\cdot G_{1}$
  • $G^{\prime}=P_{2}:1\cdot(P_{2},P_{1},P_{1})$
    $G^{\prime}=C_{3}:3\cdot(C_{3},P_{2},P_{1})$
    $G^{\prime}=G_{2}:4\cdot(G_{1},C_{3},P_{1}),3\cdot(G_{1},P_{2},P_{2})$
    $\nu_{1}=6\cdot 1+4\cdot 3+1\cdot 7=25$
  • $\nu_{2}=\binom{15}{2}=105$
@@
Since $|{\mathscr{A}}_{G_{1}}|=14$, a quadruple of potential exponents for ${\mathscr{A}}_{G_{1}}$ has to be $(1,4,4,5)_{\leq}$, or less balanced. Calculating $\operatorname{GMP}(2)$ for this tuple gives $\operatorname{GMP}(2)=69$. Since exponents that are less balanced afford a smaller $\operatorname{GMP}(2)$ and since $\operatorname{LMP}(2)=70$, we have $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so ${\mathscr{A}}_{G_{1}}$ is not free, by Lemma 2.32.
Since $|{\mathscr{A}}_{G_{1}}|=15$, choose potential balanced exponents $(1,4,5,5)_{\leq}$ for ${\mathscr{A}}_{G_{2}}$ to maximize $\operatorname{GMP}(2)$. Using this quadruple, we get $\operatorname{GMP}(2)=79$. Since exponents that are less balanced yield a smaller value for $\operatorname{GMP}(2)$ and since $\operatorname{LMP}(2)=80$, we have $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so ${\mathscr{A}}_{G_{2}}$ can’t be free, thanks to Lemma 2.32. ∎
@@
Next we modify the algorithm from Remark 4.1 to calculate $\operatorname{LMP}(2)$ when ${\mathscr{A}}_{G}$ gets equipped with a constant multiplicity $\mu=c\mathbbm{1}$ for $c\neq 1$.
@@
Algorithm to calculate $\operatorname{LMP}(2)$ for a constant multiplicity $\mu\equiv c$:
@@
    For every subgraph $G^{\prime}$ from step (1) we determine the number of triples $(G^{\prime},G[I],G[J])$, where $G[I],G[J]$ are connected induced subgraphs of $G^{\prime}$ with $N(G[I])\cup N(G[J])=N(G^{\prime})$. Let $\nu_{1}$ be the sum of the number of all such triples for all connected induced subgraphs $G^{\prime}$ from part (1).
@@
    Calculate $\nu_{2}:=\binom{|{\mathscr{A}}_{G}|}{2}$.
@@
    Since $\nu_{1}$ is the number of rank $2$ localizations with three elements, the number of rank $2$ localizations with two elements is equal to $\nu_{3}:=\nu_{2}-3\nu_{1}$, because if there is one rank $2$ localization with three elements, then the hyperplanes of this localization could have potentially formed three rank $2$ localizations with two elements.
@@
    Using Theorem 2.16 we derive the exponents of all rank $2$ localizations with three elements (which depends on the parity of $c$), while localizations of rank $2$ with two elements are free with exponents $(c,c)$. We get
[…]
@@
For $G\in\{G_{1},G_{2}\}$ and $\mu\equiv c\geq 1$ a constant multiplicity on ${\mathscr{A}}_{G}$, $({\mathscr{A}}_{G},\mu)$ is not free.
@@
The case $c=1$ is Proposition 4.3. So let $c\geq 2$. From the calculation in the example we know that: $G_{1}:$ $\nu_{1}=21,\nu_{2}=91$ and $G_{2}:$ $\nu_{1}=25,\nu_{2}=105$, which gives $\nu_{3}=28$ for $G_{1}$ and $\nu_{3}=30$ for $G_{2}$. So for $G_{1}$ we get
$\operatorname{LMP}(2)=\begin{cases}21\cdot\left(\frac{3c}{2}\right)^{2}+28c^{2}=75.25c^{2},&\text{if $c$ is even.}\\ 21\cdot\left(\frac{3c-1}{2}\cdot\frac{3c+1}{2}\right)+28c^{2}=75.25c^{2}-5.25,&\text{if $c$ is odd,}\end{cases}$
@@
$\operatorname{LMP}(2)=\begin{cases}25\cdot\left(\frac{3c}{2}\right)^{2}+30c^{2}=86.25c^{2},&\text{if $c$ is even.}\\ 25\cdot\left(\frac{3c-1}{2}\cdot\frac{3c+1}{2}\right)+30c^{2})=86.25c^{2}-6.25,&\text{if $c$ is odd.}\end{cases}$
If we assume perfectly balanced exponents for ${\mathscr{A}}_{G_{1}}$ (assuming $14c\equiv 0\mod 4$) we get the set of exponents $(\frac{14c}{4},\frac{14c}{4},\frac{14c}{4},\frac{14c}{4})_{\leq}$ which results in $\operatorname{GMP}(2)=73.5c^{2}$.
If we assume perfectly balanced exponents for ${\mathscr{A}}_{G_{2}}$ (assuming $15c\equiv 0\mod 4$) we get the set of exponents $(\frac{15c}{4},\frac{15c}{4},\frac{15c}{4},\frac{15c}{4})_{\leq}$ which results in $\operatorname{GMP}(2)=84.375c^{2}$.
Since $c\geq 2$, in any case, we have $\operatorname{LMP}(2)>\operatorname{GMP}(2)$ and thus $({\mathscr{A}}_{G_{2}},\mu)$ can’t be free, by Lemma 2.32. ∎
@@ 4.2. Constant multiplicities on a free ${\mathscr{A}}_{G}$
In this subsection, we inspect free constant multiplicities on a free ${\mathscr{A}}_{G}$. We give a complete list when $({\mathscr{A}}_{G},\mu)$ is free for $G\in\{P_{n},A_{n,k},\Delta_{n,k},C_{n}\}$ and $\mu\equiv c\geq 1$.
First we consider the case $G=P_{n}$. Terao showed in [15] that every constant multiplicity on a Coxeter arrangement is free. Since ${\mathscr{A}}_{P_{n}}$ is the Coxeter arrangement of type $A_{n}$, we derive the freeness of $({\mathscr{A}}_{P_{n}},\mu)$ for $\mu\equiv c\geq 1$.
@@
Let $G=C_{3}$ and $\mu\equiv c\geq 1$ is a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $c\in\{1,3\}$.
@@
For the forward implication, using the algorithm above, we get $\nu_{1}=1\cdot 3+3\cdot 1=6,\nu_{2}=\binom{7}{2}=21,\nu_{3}=21-3\cdot 6=3$. This gives
[…]
@@
$\operatorname{GMP}(2)=\begin{cases}16\frac{1}{3}c^{2},&\text{if }7c\equiv 0\mod 3\text{ or}\\ 16\frac{1}{3}c^{2}-\frac{1}{3},&\text{else.}\end{cases}$
For $c\geq 4$ we get that $\operatorname{LMP}(2)>\operatorname{GMP}(2)$. For $c=3$ we get equality and for $c=2$ we get $\operatorname{LMP}(2)=66$ and depending on whether the assumed exponents are $(3,5,6)_{\leq},(4,4,6)_{\leq}$, or $(4,5,5)_{\leq}$, a value of $63,64$, or $65$, respectively for $\operatorname{GMP}(2)$. So if $c\not\in\{1,3\}$ we have $\operatorname{LMP}(2)>\operatorname{GMP}(2)$ and thanks to Lemma 2.32, we deduce non-freeness.
@@
Let $G\in\{C_{n},\Delta_{n,k},A_{n,k}\}$. Suppose $\operatorname{rank}{\mathscr{A}}_{G}\geq 4$ and $\mu\equiv c\geq 1$ a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $c=1$.
@@
First let $G\in\{C_{n},\Delta_{n,k}\}$. Since $G$ either has $C_{3}$ as an induced subgraph or becomes $C_{3}$ after contracting edges, we can use Lemma 2.22 and Proposition 4.7 to focus on the cases where $c\in\{1,3\}$. If $c=1$ we know that $({\mathscr{A}}_{G},\mu)$ is free. So let $\mu\equiv 3$ and $G=C_{4}$.
Calculating $\operatorname{LMP}(2)$ we get $\nu_{1}=1\cdot(4+2)+4\cdot 2+4\cdot 1=18,\nu_{2}=\binom{13}{2}=78,\nu_{3}=78-3\cdot 18=24$ and so
@@
Let $G=A_{3,2}$. Let $\mu\equiv c$, then we get
[…]
Moreover, $\operatorname{LMP}(2)=46c^{2}$ (if $c$ is even) and $46c^{2}-3$ (if $c$ is odd). Once more we see that $\operatorname{LMP}(2)\neq\operatorname{GMP}(2)$ and so, by Theorem 2.28, $({\mathscr{A}},\mu)$ fails to be free. Since every $A_{n,k}$ with $n\geq 4$ can be transformed into $A_{3,2}$ after contracting edges, the result follows from Lemma 2.22. ∎
@@
Let $G\in\{P_{n},A_{n,k},\Delta_{n,k},C_{n}\}$ , $\mu\equiv c$ $(c>1)$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $G=P_{n}$ or $G=C_{3}$ and $c=3$.